Show me the step by step to solve the following problem.
What percent of 117 is 20.7?

Answer by Theo(13342)

you would take 20.7 and divide it by 117 and then multiply the result by 100 to get the percent.

you would get:

20.7 / 117 = .1769 rounded to 4 decimal places.

this is called the ratio.

it is also called the proportion.

.1769 * 100 = 17.69% rounded to 2 decimal places.

this is called the percent.

you show that it's a percent by including the % symbol to the right of it so that everybody knows that it's a percent and not a ratio.

ratio = percent / 100 = 17.69% / 100 = .1769.

percent = ratio * 100 = .1769 * 100 = 17.69%.

the ratio itself is 20.7 / 117 in this case.